Technical Staff Blog

Last update on .

All critical CS Department services have been restored. An interruption in external services resulted in a cascading DNS failure and overall service instability, starting mid-morning on Thursday of this week.  After repeated efforts to diagnose and workaround the issue, we reached out to CIS who stuck with us throughout today to diagnose. Ultimately, we were able to resolve the issues that were preventing our public facing DNS servers from resolving non-CS addresses.

As of 1:30am, we believe the underlying issues have been resolved and all critical services have been restored to a stable state.  Previously delayed Emails have been passing through our mail relay intermittently, thoughout the day today. And are now consistently being passed on to their destinations, without delay.

If you notice any ongoing critical system outages or otherwise are still negatively impacted as you read this, please email